# ReferencesSuppReferences

### v3

@@ Abstract
We provide precise conditions for nonparametric identification of causal effects by high-frequency event study regressions, which have been used widely in the recent macroeconomics, financial economics and political economy literatures. The high-frequency event study method regresses changes in an outcome variable on a measure of unexpected changes in a policy variable in a narrow time window around an event or a policy announcement (e.g., a 30-minute window around an FOMC announcement). We show that, contrary to popular belief, the narrow size of the window is not sufficient for identification. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other shocks (separability) and (ii) the surprise component of the news or event dominates all other shocks that are present in the event window (relative exogeneity). Technically, the latter condition requires the ratio between the variance of the policy shock and that of the other shocks to be infinite in the event window. Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity. We apply our results to Nakamura and Steinsson’s (2018a) analysis of the real economic effects of monetary policy, providing a simple empirical procedure to analyze the extent to which the standard event study estimator adequately estimates causal effects of interest.
@@ 1 Introduction
Though this approach has become central to empirical work, there are no corresponding theoretical results establishing identification of causal effects via this method, and informal identification arguments used by different authors do not always coincide. [^1] Recent examples of informal identification arguments include Bauer and Swanson (2023a) and Nakamura and Steinsson (2018a), who discussed the exogeneity and relevance of monetary policy surprises typically constructed around FOMC announcements for identifying the macroeconomic effects of monetary policy shocks. In this paper we establish precise conditions for nonparametric identification of causal effects by high-frequency event study regressions. We show that, contrary to popular belief, the narrow size of the time window the event study regression is run over is not sufficient for the identification of causal effects. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other shocks (separability) and (ii) the surprise component of the news or event dominates all other shocks that are present in the event window (relative exogeneity). Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study regression estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity (e.g., between the outcome and policy variables).
@@
Separability holds, for example, when the model is linear, which is often assumed in applied work. The key condition deserving of careful scrutiny in practice, relative exogeneity, holds when the policy shock has infinite variance while the other shocks have finite variance within the event window. More precisely, relative exogeneity holds when the ratio between the variance of the policy shock and that of the other shocks (i.e., background noise) is infinite in the event window. Thus, relative exogeneity also holds when the variance of the policy shock is finite while the variance of the background noise is vanishing. The latter shocks correspond to factors that are not specific to the announcement and may also be present in non-announcement periods. In contrast, the policy shock occurs in a lumpy manner as the unexpected part of the news quickly spreads among economic agents. It is this lumpy manner in which a disproportionate amount of policy news is revealed that can justify relative exogeneity. Even when the policy shock does not have infinite variance, which can be difficult to verify in practice, we show that the event study estimator has low bias for a weighted average of causal effects when the variance of the policy shock is large enough relative to that of other shocks in the window. In this sense, relative exogeneity can be seen as an idealized limiting case that can serve as a good approximation to the practically-relevant case of a very large variance ratio for the policy shock relative to the other shocks in the window.
@@
Relative exogeneity relates to the size of the event window. As the size of the window expands, it becomes less likely that the policy shock dominates all other shocks within the window, making it more likely that relative exogeneity provides a poor approximation. Relative exogeneity can also fail if there is information leakage about the policy news or some market anticipation of the policy change. Information leakage or frictions result in a reduction in the variance of the policy shock relative to the other shocks in the window. The finance literature has recently documented strong evidence of information leakage, informal communication and informed trading around policy announcements [see, e.g., Cieslak, Morse, and Vissing-Jorgensen (2019), Cieslak and Schrimpf (2019) and Lucca and Moench (2015)]. However, if the extent of the information leakage is small, then it will lead only to partial market anticipation and will not prevent the news from coming out in a lumpy manner at the release time, thereby allowing relative exogeneity to hold.
@@
When separability and relative exogeneity hold we show that (i) any reverse causality from the outcome variable to the policy variable does not generate bias since the reverse causality is dominated by the policy shock and (ii) common unobserved shocks correlated with the policy variable do not generate omitted variables bias since they are also dominated by the policy shock. Since the event study method regresses changes in an outcome on the unexpected changes in a policy variable at dates in the policy sample, the event study estimand is equal to the corresponding population regression coefficient. We establish that the latter identifies a weighted average of standardized marginal causal effects (MCEs) of the policy on the outcome. [^2] The MCE is the derivative of the potential outcome with respect to the policy variable. This should not be confused with the marginal treatment effect (MTE) used in microeconometrics which refers to the treatment effect of a program or intervention for those at the margin of participation [cf. Heckman and Vytlacil (2001)]. When relative exogeneity is violated, we show that the event study estimand can be decomposed into the same weighted average of MCEs and a selection bias factor. The selection bias factor is decreasing in the variance of the policy shock so that even when relative exogeneity fails, the event study estimator will not have substantial bias so long as the variance of the policy shock is relatively large.
@@
It is difficult to use statistical tests to verify relative exogeneity. Existing tests for (in)finite variance [e.g., Trapani (2016)] require a large sample of the corresponding random variable to be observed or residuals from a correctly-specified regression. Given the potential for simultaneity and omitted variables bias in high-frequency event study regressions, these requirements are not satisfied in the current context. In addition, the OLS event study estimator can still perform well when relative exogeneity holds approximately. A test for infinite variance of the policy shocks alone, even when feasible, is unable to detect when relative exogeneity provides a good enough approximation for the event study estimator to perform well in practice. Instead of trying to test for infinite variance, we introduce a simple empirical procedure that can be used for a sensitivity analysis to diagnose whether relative exogeneity is “close enough” to holding that the OLS event study estimator should be expected to have mean-squared error at least as small as that of an oracle estimator in a corresponding regression with no endogeneity. This sensitivity analysis uses information available in the relevant control sample, e.g., days when there is no FOMC meeting, in order to obtain a proxy for the variance of the shocks unrelated to the policy within the event window. This procedure requires no additional data relative to that already used by existing high-frequency event studies and can thus be applied easily. We introduce our procedure in the context of an empirical example aimed at assessing the causal effects of monetary policy news on real interest rates. The empirical results show that relative exogeneity is likely to approximately hold in the analysis of Nakamura and Steinsson (2018a) based on a 30-minute or 1-day window for the outcome and policy variable.
@@
Finally, we use our identification framework to shed light on the recent debate in the literature on some puzzling event study regression results involving the causal effect of monetary policy on Blue Chip forecasts of real GDP. We show that the estimates with signs opposite to the predictions from standard macroeconomic models likely arise because relative exogeneity fails in this type of regression. More specifically, since the Blue Chip forecast is constructed as a one-month change while the monetary policy surprise is constructed as a 30-minute change, it is highly unlikely that the variance of the policy shocks dominates the variance of the other shocks that determine the Blue Chip forecast: the latter aggregate news over a much longer time frame. Thus, any endogeneity of the policy surprise is not likely to be drowned out by the variation in the policy shock.
@@ 2.1 Event Study Design
where $X_{t}$ represents observed macroeconomic variables (or shocks) that might influence both the outcome and policy variable, $Z_{t}$ represents unobserved macroeconomic factors that also might affect the outcome and policy variable and $u_{t}$ and $e_{t}$ are serially uncorrelated shocks that are mutually uncorrelated. The parameters $\beta$ and $\alpha$ are scalars while $\theta,\,\phi,$ $\gamma_{1}$ and $\gamma_{2}$ are finite-dimensional vectors. $X_{t}$ is a vector that could include lags of $Y_{t}$ and $D_{t}$. The parameter of interest is $\beta$ which captures the response of the outcome variable to the policy action. The policy variable is endogenous, i.e., $D_{t}$ reacts to $Y_{t}$. Further, since $Z_{t}$ is not observed there is an omitted variables problem as $Z_{t}$ and $D_{t}$ may be correlated. Hence, the identification of $\beta$ requires one to overcome both simultaneity and an omitted variables problem.
The event study approach based on high-frequency observations (e.g., weekly, daily and intradaily frequencies) of $Y_{t}$ and $D_{t}$ can be used to address these identification issues. The idea is that in a narrow time window around a policy announcement or data release (e.g., FOMC announcement, U.S. employment report, GDP data release, etc.) one can extract the unexpected change (or surprise) in the policy action. The sample consists of observations of $Y_{t}$ and $D_{t}$ at the dates corresponding to the relevant policy announcement, data release or event. We call this sample the policy sample and denote it by $\mathbf{P}.$ Under the identification conditions to be discussed below, a simple OLS regression of $Y_{t}$ on $D_{t}$ over the policy sample (i.e, over all $t\in\mathbf{P})$ recovers causal effects of the policy action on the outcome.
@@ 2.2 Nonparametric Event Study Design and Identification Conditions
Assumption 2 is implied by Assumption 1, separability in $\varphi_{D}\left(Y_{t},\,Z_{t},\,e_{t},\,t\right)$ across $(Y_{t},\,Z_{t})$ and $e_{t}$ and standard invertibility requirements on $\varphi_{Y,D}\left(D_{t},\,\cdot\right)$ in $D_{t}$ and on $\varphi_{D}\left(Y_{t},\,\cdot,\,\cdot,\,\cdot\right)$ in $Y_{t}$. This holds trivially if the true data-generating process is a system of linear simultaneous equations, like (2.1) under standard rank conditions. In structural VARs, invertibility is a standard condition that is imposed to obtain the reduced-form shocks from the structural shocks [see Plagborg-Møller (2019) for a discussion about invertibility of impulse responses]. We directly impose Assumption 2 on the reduced-form rather than imposing invertibility requirements to avoid introducing further notation for defining the partial inverse of a multivariate function. Assumptions 1-2 are satisfied in the event study applications considered in empirical work since the specification for the outcome and treatment variables is typically a simultaneous linear equations model as in (2.1).
@@
In the assumed specification (2.3), $Y_{t}$ and $D_{t}$ are determined simultaneously and so are endogenous. Often endogeneity is overcome by using instrumental variables. This leads to the identification of the local average treatment effect via the instrumental variables estimand [cf. Imbens and Angrist (1994)]. In contrast, the event study approach can identify causal effects in the presence of endogeneity without the need to find instrumental variables. [^8] It is well-known that it is hard to find valid instruments in macroeconomics applications. Considering Example 1, it is difficult to find any instrument that would affect asset price returns $\left(Y_{t}\right)$ without changing short-term interest rates $\left(D_{t}\right)$ as any variable related to the macroeconomic outlook would not satisfy this criterion. Neither would variables related to corporate revenues and profits since they would likely contain information about the economic outlook and be correlated with interest rate changes. The key idea is that in a narrow time window around a particular event (or change in the policy variable) the variation in the policy variable is dominated by the variation in the policy shock. The effect of the outcome variable on the policy variable is still present within the time window but it is negligible relative to the effect of the policy shock. Similarly, the effect of the omitted variables $Z_{t}$ on $D_{t}$ is negligible relative to that of $e_{t}$. This idea is formalized by the following assumption.
@@
Assumption 4 requires that within the event window the policy shock has infinite variance (condition (i,ii)) while the other shocks have finite variance (condition (iii)) and so the policy shock dominates the changes in the policy variable in the window. Assumption 4(iii) is implied by a correspondence between the boundness of the second moments of $e_{t}$, $Z_{t}$ and $u_{t}$ with those of $D_{t}$ and $Y_{t}$. This easily holds for (2.1) since the second moments of $Y_{t}$ and $D_{t}$ depend on the second moments of $e_{t}$, $Z_{t}$ and $u_{t}$ and on the moments of their products. Altogether, Assumption 4 implies that at periods immediately surrounding a policy announcement (i.e., $t\in\mathbf{P}$) the policy shock $e_{t}$ dominates the other shock $u_{t}$ and the omitted factors $Z_{t}$. The latter involve shocks and factors that are not related to the announcement and are also present in non-announcement periods. Hence, it is reasonable to expect these shocks to have finite variance. In contrast, the policy news shock is much more pronounced in the announcement window and occurs in a lumpy manner as it is completely unexpected. Consequently, the market reacts leading to realized volatility and trading volume to significantly decline before the FOMC announcements and then jump at the announcement [see, e.g., Lucca and Moench (2015) and Hu, Pan, Wang, and Zhu (2022)]. It is this lumpy manner in which a disproportionate amount of policy news is revealed that can make the policy variable $D_{t}$ relatively exogenous. When this occurs we show that within the event window: (i) the reverse causality problem disappears (changes in the outcome variable do not affect changes in the policy variable since the latter are entirely driven by the policy shock); [^9] The same conclusion holds when observable factors are included in the analysis since these are typically low-frequency variables that are dominated by the policy shock $e_{t}$. (ii) the common unobserved factors $Z_{t}$ do not generate omitted variables bias. Note that the endogeneity problem is overcome only in the policy sample $\mathbf{P}$. In the control sample $\mathbf{C}$, defined as the collection of all $t$ such that $t\notin\mathbf{P}$ (e.g., days with no FOMC announcement), the endogeneity remains.
@@
and refer to this limiting case as relative exogeneity. [^10] More precisely, relative exogeneity holds when the ratios between the variance of the policy shock and the variances of the other shocks in the event window diverge. Thus, one may instead frame relative exogeneity as the condition for which the variance of the policy shock is finite and the variance of the background noise vanishes in the event window. The same identification results can also be shown to hold under this latter condition. The condition (2.4) makes clear one point that has been overlooked by the empirical literature. Namely, it requires that the “large” variance condition for the policy shock has to hold for all $t\in\mathbf{P}$. For example, suppose that it is satisfied only for a few announcements in the policy sample so that for the other announcements, the policy news was not a surprise for the markets. Since the variance cannot be negative, an estimate of the average variance computed in the policy sample could still be very large. This could be misleadingly interpreted as support for relative exogeneity. However, the endogeneity in most of the policy sample would not be overcome and relative exogeneity would fail. Hence, verifying that the sample variance of $D_{t}$ is large does not allow one to conclude that relative exogeneity holds.

### v4

@@ Abstract
We provide precise conditions for nonparametric identification of causal effects by high-frequency event study regressions, which have been used widely in the recent macroeconomics, financial economics and political economy literatures. The high-frequency event study method regresses changes in an outcome variable on a measure of unexpected changes in a policy variable in a narrow time window around an event or a policy announcement (e.g., a 30-minute window around an FOMC announcement). We show that, contrary to popular belief, the narrow size of the window is not sufficient for identification. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other variables (separability) and (ii) the surprise component of the news or event dominates all other variables that are present in the event window (relative exogeneity). Technically, the latter condition requires the ratio between the variance of the policy shock and that of the other variables to be infinite in the event window. Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity. We apply our results to Nakamura and Steinsson’s (2018a) analysis of the real economic effects of monetary policy, providing a simple empirical procedure to analyze the extent to which the standard event study estimator adequately estimates causal effects of interest.
@@ 1 Introduction
Though this approach has become central to empirical work, there are no corresponding theoretical results establishing identification of causal effects via this method, and informal identification arguments used by different authors do not always coincide. [^1] Recent examples of informal identification arguments include Bauer and Swanson (2023a) and Nakamura and Steinsson (2018a), who discussed the exogeneity and relevance of monetary policy surprises typically constructed around FOMC announcements for identifying the macroeconomic effects of monetary policy shocks. In this paper we establish precise conditions for nonparametric identification of causal effects by high-frequency event study regressions. We show that, contrary to popular belief, the narrow size of the time window the event study regression is run over is not sufficient for the identification of causal effects. Rather, the population regression coefficient identifies a causal estimand when (i) the effect of the policy shock on the outcome does not depend on the other variables (separability) and (ii) the surprise component of the news or event dominates all other variables that are present in the event window (relative exogeneity). Under these conditions, we establish the causal meaning of the event study estimand corresponding to the regression coefficient and the consistency and asymptotic normality of the event study regression estimator. Notably, this standard linear regression estimator is robust to general forms of nonlinearity (e.g., between the outcome and policy variables).
@@
Separability holds, for example, when the model is linear, which is often assumed in applied work. The key condition deserving of careful scrutiny in practice, relative exogeneity, holds when the policy shock has infinite variance while the other variables have finite variance within the event window. More precisely, relative exogeneity holds when the ratio between the variance of the policy shock and that of the other variables (i.e., background noise) is infinite in the event window. Thus, relative exogeneity also holds when the variance of the policy shock is finite while the variance of the background noise is vanishing. The latter variables correspond to factors that are not specific to the announcement and may also be present in non-announcement periods. In contrast, the policy shock occurs in a lumpy manner as the unexpected part of the news quickly spreads among economic agents. It is this lumpy manner in which a disproportionate amount of policy news is revealed that can justify relative exogeneity. Even when the policy shock does not have infinite variance, which can be difficult to verify in practice, we show that the event study estimator has low bias for a weighted average of causal effects when the variance of the policy shock is large enough relative to that of other variables in the window. In this sense, relative exogeneity can be seen as an idealized limiting case that can serve as a good approximation to the practically-relevant case of a very large variance ratio for the policy shock relative to the other variables in the window.
@@
Relative exogeneity relates to the size of the event window. As the size of the window expands, it becomes less likely that the policy shock dominates all other variables within the window, making it more likely that relative exogeneity provides a poor approximation. Relative exogeneity can also fail if there is information leakage about the policy news or some market anticipation of the policy change. Information leakage or frictions result in a reduction in the variance of the policy shock relative to the other variables in the window. The finance literature has recently documented strong evidence of information leakage, informal communication and informed trading around policy announcements [see, e.g., Cieslak, Morse, and Vissing-Jorgensen (2019), Cieslak and Schrimpf (2019) and Lucca and Moench (2015)]. However, if the extent of the information leakage is small, then it will lead only to partial market anticipation and will not prevent the news from coming out in a lumpy manner at the release time, thereby allowing relative exogeneity to hold.
@@
When separability and relative exogeneity hold we show that (i) any reverse causality from the outcome variable to the policy variable does not generate bias since the reverse causality is dominated by the policy shock and (ii) common unobserved factors correlated with the policy variable do not generate omitted variables bias since they are also dominated by the policy shock. Since the event study method regresses changes in an outcome on the unexpected changes in a policy variable at dates in the policy sample, the event study estimand is equal to the corresponding population regression coefficient. We establish that the latter identifies a weighted average of standardized marginal causal effects (MCEs) of the policy on the outcome. [^2] The MCE is the derivative of the potential outcome with respect to the policy variable. This should not be confused with the marginal treatment effect (MTE) used in microeconometrics which refers to the treatment effect of a program or intervention for those at the margin of participation [cf. Heckman and Vytlacil (2001)]. When relative exogeneity is violated, we show that the event study estimand can be decomposed into the same weighted average of MCEs and a selection bias factor. The selection bias factor is decreasing in the variance of the policy shock so that even when relative exogeneity fails, the event study estimator will not have substantial bias so long as the variance of the policy shock is relatively large.
@@
It is difficult to use statistical tests to verify relative exogeneity. Existing tests for (in)finite variance [e.g., Trapani (2016)] require a large sample of the corresponding random variable to be observed or residuals from a correctly-specified regression. Given the potential for simultaneity and omitted variables bias in high-frequency event study regressions, these requirements are not satisfied in the current context. In addition, the OLS event study estimator can still perform well when relative exogeneity holds approximately. A test for infinite variance of the policy shocks alone, even when feasible, is unable to detect when relative exogeneity provides a good enough approximation for the event study estimator to perform well in practice. Instead of trying to test for infinite variance, we introduce a simple empirical procedure that can be used for a sensitivity analysis to diagnose whether relative exogeneity is “close enough” to holding that the OLS event study estimator should be expected to have mean-squared error at least as small as that of an oracle estimator in a corresponding regression with no endogeneity. This sensitivity analysis uses information available in the relevant control sample, e.g., days when there is no FOMC meeting, in order to obtain a proxy for the variance of the variables that are not specific to the policy within the event window. This procedure requires no additional data relative to that already used by existing high-frequency event studies and can thus be applied easily. We introduce our procedure in the context of an empirical example aimed at assessing the causal effects of monetary policy news on real interest rates. The empirical results show that relative exogeneity is likely to approximately hold in the analysis of Nakamura and Steinsson (2018a) based on a 30-minute or 1-day window for the outcome and policy variable.
@@
Finally, we use our identification framework to shed light on the recent debate in the literature on some puzzling event study regression results involving the causal effect of monetary policy on Blue Chip forecasts of real GDP. We show that the estimates with signs opposite to the predictions from standard macroeconomic models likely arise because relative exogeneity fails in this type of regression. Since the Blue Chip forecast revision is constructed as a one-month change while the monetary policy surprise is constructed as a 30-minute change, it is highly unlikely that the variance of the policy shocks dominates the variance of the other variables that determine the Blue Chip forecast: the latter aggregate news over a much longer time frame. Thus, any endogeneity of the policy surprise is not likely to be drowned out by the variation in the policy shock.
@@ 2.1 Event Study Design
where $X_{t}$ represents observed macroeconomic variables that might influence both the outcome and policy variable, $Z_{t}$ represents unobserved macroeconomic factors that also might affect the outcome and policy variable and $u_{t}$ and $e_{t}$ are serially uncorrelated shocks that are mutually uncorrelated. The parameters $\beta$ and $\alpha$ are scalars while $\theta,\,\phi,$ $\gamma_{1}$ and $\gamma_{2}$ are finite-dimensional vectors. $X_{t}$ is a vector that could include lags of $Y_{t}$ and $D_{t}$. The parameter of interest is $\beta$ which captures the response of the outcome variable to the policy action. The policy variable is endogenous, i.e., $D_{t}$ reacts to $Y_{t}$. Further, since $Z_{t}$ is not observed there is an omitted variables problem as $Z_{t}$ and $D_{t}$ may be correlated. Hence, the identification of $\beta$ requires one to overcome both simultaneity and an omitted variables problem.
The event study approach based on high-frequency observations (e.g., weekly, daily and intradaily frequencies) of $Y_{t}$ and $D_{t}$ can be used to address these identification issues. The idea is that in a narrow time window around a policy announcement or data release (e.g., FOMC announcement, U.S. employment report, GDP data release, etc.) one can extract the unexpected change (or surprise) in the policy action to form $D_{t}$. The sample consists of observations of $Y_{t}$ and $D_{t}$ at the dates corresponding to the relevant policy announcement, data release or event. We call this sample the policy sample and denote it by $\mathbf{P}.$ Under the identification conditions to be discussed below, a simple OLS regression of $Y_{t}$ on $D_{t}$ over the policy sample (i.e, over all $t\in\mathbf{P})$ recovers causal effects of the policy action on the outcome.
@@ 2.2 Nonparametric Event Study Design and Identification Conditions
Assumption 2 is implied by Assumption 1, separability in $\varphi_{D}\left(Y_{t},\,Z_{t},\,e_{t},\,t\right)$ across $(Y_{t},\,Z_{t})$ and $e_{t}$ and standard invertibility requirements on the system of equations. This holds trivially if the true data-generating process is a system of linear simultaneous equations, like (2.1) under standard rank conditions. In structural VARs, invertibility is a standard condition that is imposed to obtain the reduced-form shocks from the structural shocks [see Plagborg-Møller (2019) for a discussion about invertibility of impulse responses]. We directly impose Assumption 2 on the reduced-form rather than imposing invertibility requirements to avoid introducing further notation for defining the partial inverse of a multivariate function. Assumptions 1-2 are satisfied in the event study applications considered in empirical work since the specification for the outcome and treatment variables is typically a simultaneous linear equations model as in (2.1).
@@
In the assumed specification (2.3), $Y_{t}$ and $D_{t}$ are determined simultaneously and so are endogenous. Often endogeneity is overcome by using instrumental variables. This leads to the identification of the local average treatment effect via the instrumental variables estimand [cf. Imbens and Angrist (1994)]. In contrast, the event study approach can identify causal effects in the presence of endogeneity without the need to find instrumental variables. [^8] It is well-known that it is hard to find valid instruments in macroeconomics applications. Considering Example 1, it is difficult to find any instrument that would affect asset price returns $\left(Y_{t}\right)$ without changing short-term interest rates $\left(D_{t}\right)$ as any variable related to the macroeconomic outlook would not satisfy this criterion. Neither would variables related to corporate revenues and profits since they would likely contain information about the economic outlook and be correlated with interest rate changes. The key idea is that in a narrow time window around a particular event (or change in policy) the variation in the policy variable is dominated by the variation in the policy shock. The effect of the outcome variable on the policy variable is still present within the time window but it is negligible relative to the effect of the policy shock. Similarly, the effect of the omitted variables $Z_{t}$ on $D_{t}$ is negligible relative to that of $e_{t}$. This idea is formalized by the following assumption.
@@
Assumption 4 requires that within the event window the policy shock has infinite variance (condition (i,ii)) while the other variables have finite variance (condition (iii)) and so the policy shock dominates the changes in the policy variable in the window. Assumption 4(iii) is implied by a correspondence between the boundness of the second moments of $e_{t}$, $Z_{t}$ and $u_{t}$ with those of $D_{t}$ and $Y_{t}$. This easily holds for (2.1) since the second moments of $Y_{t}$ and $D_{t}$ depend on the second moments of $e_{t}$, $Z_{t}$ and $u_{t}$ and on the moments of their products. Altogether, Assumption 4 implies that at periods immediately surrounding a policy announcement (i.e., $t\in\mathbf{P}$) the policy shock $e_{t}$ dominates the other shock $u_{t}$ and the omitted factors $Z_{t}$. The latter involve shocks and factors that are not related to the announcement and are also present in non-announcement periods. Hence, it is reasonable to expect these variables to have finite variance. In contrast, the policy news shock is much more pronounced in the announcement window and occurs in a lumpy manner as it is completely unexpected. Consequently, the market reacts leading to realized volatility and trading volume to significantly decline before the announcement and then jump at the announcement [see, e.g., Lucca and Moench (2015) and Hu, Pan, Wang, and Zhu (2022)]. It is this lumpy manner in which a disproportionate amount of policy news is revealed that can make the policy variable $D_{t}$ relatively exogenous. When this occurs we show that within the event window: (i) the reverse causality problem disappears (changes in the outcome variable do not affect changes in the policy variable since the latter are entirely driven by the policy shock); [^9] The same conclusion holds when observable factors are included in the analysis since these are typically low-frequency variables that are dominated by the policy shock $e_{t}$. (ii) the common unobserved factors $Z_{t}$ do not generate omitted variables bias. Note that the endogeneity problem is overcome only in the policy sample $\mathbf{P}$. In the control sample $\mathbf{C}$, defined as the collection of all $t$ such that $t\notin\mathbf{P}$ (e.g., days with no FOMC announcement), the endogeneity remains.
@@
and refer to this limiting case as relative exogeneity. [^10] More precisely, relative exogeneity holds when the ratios between the variance of the policy shock and the variances of the other variables in the event window diverge. Thus, one may instead frame relative exogeneity as the condition for which the variance of the policy shock is finite and the variance of the background noise vanishes in the event window. The same identification results can also be shown to hold under this latter condition. The condition (2.4) makes clear one point that has been overlooked by the empirical literature. Namely, it requires that the “large” variance condition for the policy shock has to hold for all $t\in\mathbf{P}$. For example, suppose that it is satisfied only for a few announcements in the policy sample. Since the variance cannot be negative, an estimate of the average variance computed in the policy sample could still be very large. This could be misleadingly interpreted as support for relative exogeneity. However, the endogeneity in most of the policy sample would not be overcome and relative exogeneity would fail. Hence, verifying that the sample variance of $D_{t}$ is large does not allow one to conclude that relative exogeneity holds.